OpenStudy (anonymous):

Solve for x algebraically in 7^(-x+2) = 0.005, rounding your answers to three signicant digits. A) 4.32 B) 4.72 C) 5.12 D) 5.52 E) 5.92

OpenStudy (anonymous):

You'll need to use logs for this.

OpenStudy (anonymous):

Do I use 7ln(-x+2) = ln 0.005 ???

OpenStudy (anonymous):

You do take the log of both sides, but I would suggest using log base 10 and what you have there isn't quite correct even if you were going to use the natural log.

OpenStudy (anonymous):

Remember that \[log(a^b) = b*log(a)\]